# JiSAM: Alleviate Labeling Burden and Corner Case Problems in Autonomous Driving via Minimal Real-World Data

### v2

@@ Abstract
Deep-learning-based autonomous driving (AD) perception introduces a promising picture for safe and environment-friendly transportation. However, the over-reliance on real labeled data in LiDAR perception limits the scale of on-road attempts. 3D real world data is notoriously time-and-energy-consuming to annotate and lacks corner cases like rare traffic participants. On the contrary, in simulators like CARLA, generating labeled LiDAR point clouds with corner cases is a piece of cake. However, introducing synthetic point clouds to improve real perception is non-trivial. This stems from two challenges: 1) sample efficiency of simulation datasets 2) simulation-to-real gaps. To overcome both challenges, we propose a plug-and-play method called JiSAM , shorthand for Jittering augmentation, domain-aware backbone and memory-based Sectorized AlignMent. In extensive experiments conducted on the famous AD dataset NuScenes, we demonstrate that, with SOTA 3D object detector, JiSAM  is able to utilize the simulation data and only labels on $2.5\%$ available real data to achieve comparable performance to models trained on all real data. Additionally, JiSAM  achieves more than 15 mAPs on the objects not labeled in the real training set. We will release models and codes.
@@ 3.1 Notations and Problem Formulation
Notations. To begin with, we denote LiDAR point cloud $\mathbf{P}\in\mathbb{R}^{N\times(3+d)}$ as the concatenation of the $xyz$-coordinate $\mathbf{C}\in\mathbb{R}^{N\times 3}$ and the point features $\mathbf{F}\in\mathbb{R}^{N\times d}$, which leads to $\mathbf{P}=[\mathbf{C},\mathbf{F}]$. Here $N$ denotes the number of points (or voxels) and $d$ is the number of point feature channels, which varies among different datasets. For example, $d=2$ in NuScenes dataset [3] representing the intensity and collected timestamp of each point. Moreover, as the point intensity from CARLA simulator is computed via a linear function of xyz-coordinate, there is no more useful information from this only feature in simulation dataset and $d=0$ for all the simulation datasets. Each LiDAR sample is paired with the bounding box labels $\mathbf{Y}\in\mathbb{R}^{M\times 10}$ for 3D object detection, where $M$ is the number of the surrounding objects and each 3D bounding box contains 10 scalars indicating category, xyz-location, xyz-size (length, width and height), orientation with respect to z-axis ($\alpha$) and speeds along x and y axis respectively. Thus each data sample $\mathbf{D}=[\mathbf{P},\mathbf{Y}]$ is consisted by the LiDAR point cloud and the labels. To indicate data samples from different domains, we use the subscripts ‘real’ and ‘sim’. For instance, real-world and synthetic point clouds are represented by $\mathbf{P}_{\text{real}}=[\mathbf{C}_{\text{real}},\mathbf{F}_{\text{real}}]\in\mathbb{R}^{N_{\text{real}}\times(3+d)}$ and $\mathbf{P}_{\text{sim}}=[\mathbf{C}_{\text{sim}}]\in\mathbb{R}^{N_{\text{sim}}\times 3}$ respectively.
@@
$\hat{\mathbf{Y}}=f^{\text{DET}}(\hat{\mathbf{B}})$ (3)
Problem Formulation. For a real-world dataset $\mathcal{D}_{real}$, we first simulate a similar LiDAR configuration in the CARLA simulator and generate large amount of synthetic data with bounding box labels. Then we build a joint dataset $\mathcal{D}=\{\hat{\mathcal{D}}_{\text{real}},\mathcal{D}_{\text{sim}}\}$ with fewer real training samples $\hat{\mathcal{D}}_{\text{real}}=\{[\mathbf{P}_{\text{real}},\mathbf{Y}_{\text{real}}]\}$ and large amount of synthetic data $\mathcal{D}_{\text{sim}}=\{[\mathbf{P}_{\text{sim}},\mathbf{Y}_{\text{sim}}]\}$. In our experiments, $|\hat{\mathcal{D}}_{\text{real}}|\approx 7,000$ and $|\mathcal{D}_{\text{sim}}|\approx 80,000$, where ||$\cdot$|| ⋅ | indicates the size of a set. Then we train current SOTA 3D detection models on $\mathcal{D}$ and tested on the whole validation set in $\mathcal{D}_{real}$.
@@
$\mathcal{L}=\mathcal{L}^{\text{DET}}(\hat{\mathbf{Y}},\mathbf{Y})+\lambda\mathcal{L}^{\text{SMA}}(\hat{\mathbf{B}})$ (4)
@@ 3.2 Noise Augmentation
[…]
@@
[…]
@@
{$r&=\sqrt{x^{2}+y^{2}+z^{2}}\\ \theta&=\arccos{\frac{z}{r}}\\ \varphi&=\operatorname{sgn}(y)\arccos{\frac{x}{\sqrt{x^{2}+y^{2}}}}\end{aligned}\right.\ \ ,\ \ \operatorname{sgn}(y)=\left\{\begin{aligned} -1\ \ &\text{if}\ y<0,\\ 0\ \ \ &\text{if}\ y=0,\\ 1\ \ \ &\text{if}\ y>0.\end{aligned}\right.$ (7)
@@ 3.4 Alignment Losses
In practice, we find that conducting the memory-based sectorized alignment loss in a bi-directional way introduce more improvement, which means that we separately maintain memory banks for both real $\mathbf{M}_{\text{real}}$ and simulation objects $\mathbf{M}_{\text{sim}}$ and compute the alignment loss with $<\mathbf{M}_{\text{real}},\mathbf{O}_{\text{sim}}>$ and $<\mathbf{M}_{\text{sim}},\mathbf{O}_{\text{real}}>$ pairs. Note that JiSAM does not construct real memory for categories not in real data but if they exists in synthetic data, the simulation bank is created. Categories available in the memory banks are involved in computing $\mathcal{L}^{\text{SMA}}$.

### v3

# arXiv:2503.08422v3[cs.CV] 27 Feb 2026

@@ Abstract
Deep-learning-based autonomous driving (AD) perception introduces a promising picture for safe and environment-friendly transportation. However, the over-reliance on real labeled data in LiDAR perception limits the scale of on-road attempts. 3D real world data is notoriously time-and-energy-consuming to annotate and lacks corner cases like rare traffic participants. On the contrary, in simulators like CARLA, generating labeled LiDAR point clouds with corner cases is a piece of cake. However, introducing synthetic point clouds to improve real perception is non-trivial. This stems from two challenges: 1) sample efficiency of simulation datasets 2) simulation-to-real gaps. To overcome both challenges, we propose a plug-and-play method called JiSAM , shorthand for Jittering augmentation, domain-aware backbone and memory-based Sectorized AlignMent. In extensive experiments conducted on the famous AD dataset NuScenes, we demonstrate that, with SOTA 3D object detector, JiSAM is able to utilize the simulation data and only labels on $2.5\%$ available real data to achieve comparable performance to models trained on all real data. Additionally, JiSAM achieves more than 15 mAPs on the objects not labeled in the real training set. Codes and models are available here
@@ 3.1 Notations and Problem Formulation
Notations. To begin with, we denote LiDAR point cloud $\mathbf{P}\in\mathbb{R}^{N\times(3+d)}$ as the concatenation of the xyzxyz-coordinate $\mathbf{C}\in\mathbb{R}^{N\times 3}$ and the point features $\mathbf{F}\in\mathbb{R}^{N\times d}$, which leads to $\mathbf{P}=[\mathbf{C},\mathbf{F}]$. Here $N$ denotes the number of points (or voxels) and $d$ is the number of point feature channels, which varies among different datasets. For example, $d=2$ in NuScenes dataset [3] representing the intensity and collected timestamp of each point. Moreover, as the point intensity from CARLA simulator is computed via a linear function of xyz-coordinate, there is no more useful information from this only feature in simulation dataset and $d=0$ for all the simulation datasets. Each LiDAR sample is paired with the bounding box labels $\mathbf{Y}\in\mathbb{R}^{M\times 10}$ for 3D object detection, where $M$ is the number of the surrounding objects and each 3D bounding box contains 10 scalars indicating category, xyz-location, xyz-size (length, width and height), orientation with respect to z-axis ($\alpha$) and speeds along x and y axis respectively. Thus each data sample $\mathbf{D}=[\mathbf{P},\mathbf{Y}]$ is consisted by the LiDAR point cloud and the labels. To indicate data samples from different domains, we use the subscripts ‘real’ and ‘sim’. For instance, real-world and synthetic point clouds are represented by $\mathbf{P}_{\text{real}}=[\mathbf{C}_{\text{real}},\mathbf{F}_{\text{real}}]\in\mathbb{R}^{N_{\text{real}}\times(3+d)}$ and $\mathbf{P}_{\text{sim}}=[\mathbf{C}_{\text{sim}}]\in\mathbb{R}^{N_{\text{sim}}\times 3}$ respectively.
@@
[…]
Problem Formulation. For a real-world dataset $\mathcal{D}_{real}$, we first simulate a similar LiDAR configuration in the CARLA simulator and generate large amount of synthetic data with bounding box labels. Then we build a joint dataset $\mathcal{D}=\{\hat{\mathcal{D}}_{\text{real}},\mathcal{D}_{\text{sim}}\}$ with fewer real training samples $\hat{\mathcal{D}}_{\text{real}}=\{[\mathbf{P}_{\text{real}},\mathbf{Y}_{\text{real}}]\}$ and large amount of synthetic data $\mathcal{D}_{\text{sim}}=\{[\mathbf{P}_{\text{sim}},\mathbf{Y}_{\text{sim}}]\}$. In our experiments, $|\hat{\mathcal{D}}_{\text{real}}|\approx 7,000$ and $|\mathcal{D}_{\text{sim}}|\approx 80,000$, where $|\cdot|$ indicates the size of a set. Then we train current SOTA 3D detection models on $\mathcal{D}$ and tested on the whole validation set in $\mathcal{D}_{real}$.
@@
[…]
@@ 3.2 Noise Augmentation
$\left\{\begin{aligned} \dot{r}&=r+\delta_{r},&\delta_{r}\sim N(0,\Delta_{r})\\ \dot{\theta}&=\theta+\delta_{\theta},&\delta_{\theta}\sim N(0,\Delta_{\theta})\\ \dot{\varphi}&=\varphi+\delta_{\varphi},&\delta_{\varphi}\sim N(0,\Delta_{\varphi})\end{aligned}\right$. (5)
@@
$\left\{\begin{aligned} x&=\dot{r}\sin{\dot{\theta}}\cos{\dot{\varphi}}\\ y&=\dot{r}\sin{\dot{\theta}}\sin{\dot{\varphi}}\\ z&=\dot{r}\cos{\dot{\theta}}\end{aligned}\right$. (6)
@@
$\left\{\begin{aligned} r&=\sqrt{x^{2}+y^{2}+z^{2}}\\ \theta&=\arccos{\frac{z}{r}}\\ \varphi&=\operatorname{sgn}(y)\arccos{\frac{x}{\sqrt{x^{2}+y^{2}}}}\end{aligned}\right.\ \ ,\ \ \operatorname{sgn}(y)=\left\{\begin{aligned} -1\ \ &\text{if}\ y<0,\\ 0\ \ \ &\text{if}\ y=0,\\ 1\ \ \ &\text{if}\ y>0.\end{aligned}\right$. (7)
@@ 3.4 Alignment Losses
In practice, we find that conducting the memory-based sectorized alignment loss in a bi-directional way introduce more improvement, which means that we separately maintain memory banks for both real $\mathbf{M}_{\text{real}}$ and simulation objects $\mathbf{M}_{\text{sim}}$ and compute the alignment loss with $\mathbf{M}_{\text{real}},\mathbf{O}_{\text{sim}}>$ and $\mathbf{M}_{\text{sim}},\mathbf{O}_{\text{real}}>$ pairs. Note that JiSAM does not construct real memory for categories not in real data but if they exists in synthetic data, the simulation bank is created. Categories available in the memory banks are involved in computing $\mathcal{L}^{\text{SMA}}$.